Pearl Corp. is expected to have an EBIT of $2,700,000 next year. Depreciation, the increase in net working capital, and capital spending are expected to be $160,000, $120,000, and $160,000, respectively. All are expected to grow at 17 percent per year for four years. The company currently has $14,000,000 in debt and 1,200,000 shares outstanding. After Year 5, the adjusted cash flow from assets is expected to grow at 3.5 percent indefinitely. The company’s WACC is 9.3 percent and the tax rate is 24 percent. What is the price per share of the company's stock? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

HI

Net Income =EBIT *[1-Tax]

                 = 2,700,000*[1-.24]

                  = $2,052,000

Free cash flow for year 1 =Net Income + Depreciation -increase in working capital-capital spending

              = 2,052,000+160,000-120,000-160,000

            = $1,932,000

FCF2=1932000(1+.17]= $2,260,440

FCF3= 2260440[1+.17]= 2,644,714.80

FCF4=2644714.80[1+.17]= $3,094,316.32

FCF5= 3094316.32*[1+.17]= $3,620,350.09

Terminal value at year 5= FCF5(1+g]/(WACC-g]

               = 3620350.09*(1+.035)/(.093-.035)

                =$64,604,523.15

Now we will calculate the sum of present value of these free cash flows which will be value of company

Value of company = 1932000/(1+9.3%) + 2260440/(1+9.3%)^2 + 2644714.80/(1+9.3%)^3 + 3094316.32/(1+9.3%)^4

+ 3620350.09/(1+9.3%)^5 + 64604523.15/(1+9.3%)^5

= $51,589,597.08

Value of equity = Value of company - Value of debt

= 51589597.08 - 14,000,000

= $37,589,597.08

Per Share Value = 37589597.08/1,200,000

= $31.32

Thanks